40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Mar 28 03:37:07 2013, Zach, Optics, Configuration, Test setup input optics progress OMC_test_setup_IO.jpg2013-03-27_05.48.20.jpg2013-03-28_02.29.36.jpg
    Reply  Fri Mar 29 08:55:00 2013, Zach, Optics, Configuration, Beam launched into fiber 
       Reply  Thu Apr 4 23:44:52 2013, Koji, Optics, Configuration, Beam launched into fiber 
          Reply  Fri Apr 5 18:18:36 2013, Zach, Optics, Configuration, AOM probably broken Isomet_1250C_zin_4_5_13.png
Message ID: 99     Entry time: Fri Apr 5 18:18:36 2013     In reply to: 97
Author: Zach 
Type: Optics 
Category: Configuration 
Subject: AOM probably broken 

Quote:

Then, I started to check the AOM path. I noticed that the 1st (or -1st) order beam is very weak.
The deflection efficiency is ~0.1%. Something is wrong.
I checked the driver. The driver's coupler output (1:10) show the amplitude ~1V. (good)
I check the main output by reducing the offset. When the coupler output is 100mV, the main output was 1V. (good)
So is the AOM itself broken???

As Koji noticed that the AOM efficiency was very low, I figured I would try looking at it with a fresh set of eyes. The end result is that I have to agree that the AOM appears to be broken.

First, I measured the input impedance of the AOM using the AG4395A with the impedance test kit (after calibrating). The plot is below. The spec sheet says the center frequency is 200 MHz, at which Zin should be ~50 ohms. It crosses 50 ohms somewhere near 235 MHz, which may be reasonable given that the LC circuit can be tuned by hand. However, it does surprise me that the impedance varies so much over the specified RF range of ±50 MHz. Maybe this is an indication that something is bad.

Isomet_1250C_zin_4_5_13.png

I removed the cover of the modulator (which I think Koji did, as well) and all the connections looked as I imagine they should---i.e., there was nothing obviously broken, physically.

I then tried my hand at realigning the AOM from scratch by removing and replacing it. I was not able to get better than 0.15%, which is roughly what Koji got.

So, perhaps our best course of action is to decide what we expect the Zin spectrum to look like, and whether that agrees with the above measurement.

ELOG V3.1.3-